prettygirl15:

What is (f–g)(x)? f(x)=2x+4 g(x)=–3x

kittybasil:

This (click the blue text) might help you on the concept involved with this problem.

kittybasil:

kittybasil wrote:
This (click the blue text) might help you on the concept involved with this problem.
Anyway. @prettygirl15 based on the link I shared, do you get what should be done here?

Florisalreadytaken:

k so based off the info we just got, OR what we already know lets solve this! \[ (f–g)(x) \Rightarrow f(x) - g(x) \] look at what we're given: \[ f(x)=2x+4 \] \[ g(x)=–3x \] that said, just plug those in the formula above: \[ 2x+4=–3x \rightarrow 5x=-4 \rightarrow x= \frac{-4}{5} \Rightarrow x= -\frac{4}{5} \] yes?

prettygirl15:

thx

Florisalreadytaken:

always!